what is the square root of -i argument of -i

what is the


square root of -i


argument of -i

Grade:11

4 Answers

Chetan Mandayam Nayakar
11 years ago

sqrt(-i)=e-i((pi/2)+(2N(pi)))

where N is ANY integer

Akash Kumar Dutta
11 years ago

find sq root by the simple step... root (-i)=a+ib...............
and arg of -i= -pi/2 or 5pi/2

sandy sandy
11 years ago

arg is 270 ad 1is the root

Harsh Gupta
10 years ago

Argument if -i is -π/2.